OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-4, 6) and (2, -6). Please show all of your work.

myininaya (myininaya):

\[m=\frac{y_2-y_1}{x_2-x_1}\] use one of the points \[(x_1,y_1)\] and plug into y=mx+b \[y_1=mx_1+b\] solve for b to find y-intercept

OpenStudy (anonymous):

6+6/-4-2=12/-6=-2 so m=-2 6=-2(-4)+b 6=8+b b=-2 so y=-2x-2

OpenStudy (anonymous):

Thanks a lot, for explaining.